# CharacterFactory: Sampling Consistent Characters with GANs for Diffusion Models

### v1

@@ 4.3. Comparison with Consistent-T2I Methods
In addition, we compare the most related methods The Chosen One (avrahami2023chosen, ) and ConsiStory (tewel2024training, ) with the content provided in their papers. These two methods are also designed for consistent character generation, but have not released the codes yet.
@@ 4.5. Interpolation Property of IDE-GAN
The interpolation property of GANs is that interpolations between different randomly sampled latent codes in latent space can produce semantically smooth variations in image space (sainburg2018generative, ). To evaluate whether our IDE-GAN carries this property, we randomly sample $z_{1}$ and $z_{2}$, and perform linear interpolation as shown in Figure 6. IDE-GAN uses the interpolated latent codes to generate corresponding pseudo identity embeddings, respectively. Since the output space of IDE-GAN is embeddings instead of images, it cannot directly visualize the variations like traditional GANs (sainburg2018generative, ; karras2019style, ) in image space. So we insert these pseudo identity embeddings into Stable Diffusion to generate the corresponding images via the pipeline in Figure 2(b). As shown in Figure 6, CharacterFactory can produce continuous identity variations with the interpolations between different latent codes. And the interpolated latent codes (e.g., $0.5z_{1}+0.5z_{2}$) can be chosen for further identity-consistent generation. It demonstrates that our IDE-GAN has satisfactory interpolation property and can be seamlessly integrated with Stable Diffusion.
@@ 4.6. Applications
Virtual Humans in Image/Video/3D Generation. Currently, virtual human generation mainly includes 2D/3D facial reconstruction, talking-head generation and body/human movements (zhen2023human, ), which typically rely on pre-existing images and lack scenario diversity and editability. And CharacterFactory can create new characters end-to-end and conduct identity-consistent virtual human image generation. In addition, since the pretrained Stable Diffusion 2.1 is fixed and the generated pseudo identity embeddings can be inserted into CLIP text transformer naturally, our method can collaborate with the SD-based plug-and-play modules. As shown in Figure 1, we integrate CharacterFactory with ControlNet-OpenPose (zhang2023adding, ; cao2017realtime, ), ModelScopeT2V (wang2023modelscope, ) and LucidDreamer (liang2023luciddreamer, ) to implement identity-consistent virtual human image/video/3D generation.
Identity-Consistent Dateset Construction. Some human-centric subject-driven generation methods (li2023photomaker, ; chen2023dreamidentity, ) construct large-scale celeb datasets for training. PhotoMaker (li2023photomaker, ) crawls celeb photos from the Internet and DreamIdentity (chen2023dreamidentity, ) uses text prompts containing celeb names to drive Stable Diffusion to generate celeb images. Their constructed data includes only celebs, leading to a limited number of identities. Notably, the proposed CharacterFactory can use diverse text prompts to generate identity-consistent images infinitely for dataset construction. Furthermore, collaboration with the mentioned SD-based plug-and-play modules can construct identity-consistent video/3D datasets.
@@ Appendix A More Results
In addition, we provide identity-consistent Image/Video/3D generation results as shown in Figure 9. Seamless integration with ControlNet (zhang2023adding, ), ModelScopeT2V (wang2023modelscope, ) and LucidDreamer (liang2023luciddreamer, ) empowers identity-consistent virtual human image/video/3D generation.
Comparison with Two-Stage Workflows. As mentioned in Section 1 of the main paper, a feasible way to create a new character is that a text-to-image model is used to create a new character’s image and then subject-driven methods are adopted to produce images based on the pre-generated image. We conduct experiments with the competing methods under this setting as shown in Figure 10 and Table 4. Obviously, the two-stage workflows directly lead to an unacceptable degradation of competing methods on image quality qualitatively and quantitatively. On the other hand, two-stage workflows consume more time for creating identity-consistent characters. In comparison, our end-to-end framework ensures both strong consistent generation performance and faster inference workflow.
Figure 10. Qualitative comparisons with two-stage workflows using five baselines (denoted with $\dagger$) for creating consistent characters. In comparison, the proposed CharacterFactory can generate high-quality identity-consistent character images with diverse layouts while conforming to the given text prompts.
Table 4. Quantitative comparisons with two-stage workflows using five baselines (denoted with $\dagger$) for creating consistent characters. We define the speed as the time it takes to create a new consistent character on a single NVIDIA A100 GPU. Obviously, CharacterFactory obtains superior performance on identity consistency, editability, trusted face diversity, image quality and speed, which are consistent with the qualitative comparisons.
[…]
Textual Inversion (gal2022image, ) 0.647 0.295 0.274 0.392 0.078 47.94 3200
DreamBooth (ruiz2023dreambooth, ) 0.681 0.443 0.287 0.339 0.073 62.66 1500
IP-Adapter (ye2023ip, ) 0.853 0.447 0.227 0.192 0.096 95.25 7
Celeb-Basis (yuan2023inserting, ) 0.667 0.369 0.273 0.378 0.101 56.43 480
PhotoMaker (li2023photomaker, ) 0.694 0.451 0.301 0.331 0.138 53.37 10
CharacterFactory 0.764 0.498 0.332 0.333 0.140 22.58 3
@@ Appendix B More Implementation Details
The Dimension of Latent Code $z$. Similar to traditional GAN models (karras2019style, ; fu2022stylegan, ), the input of our Identity-Embedding GAN (IDE-GAN) is randomly sampled from a latent space, i.e., $z\in\mathcal{N}(0,\mathbf{I})$. Differently, our objective is to learn a mapping from the latent space to the celeb embedding space, i.e., $G:z\rightarrow v$. The CLIP word embedding’s dimension of Stable Diffusion 2.1 is 1024, so the dimension of IDE-GAN’s output $[v_{1}^{*},v_{2}^{*}]$ is $[2,1024]$. Based on this, we conduct experiments for the dimension of $z$, as shown in Table 5. Overall, IDE-GAN can generate consistent pseudo identity conforming to the given text prompts with different dimensions of $z$. However, we empirically find that a larger dimension of $z$ introduces more trainable parameters, leading to harder convergence and degraded identity consistency. A smaller $z$ represents narrower input space, resulting in poor identity diversity (this metric is designed in Section 4.4 for the diversity of generated identities). Therefore, we choose the setting of “Dim(z)=64” to trade off identity consistency and identity diversity, which performs well across the board.
@@
More Details of IDE-GAN. IDE-GAN is made up of MLPs, where the generator is $64-2048-2048$ MLP architecture and the discriminator is $2048-512-256-1$ MLP architecture. In addition to the ablation experiments in Section 4.4 of the main paper, we further explore the effect of AdaIN (wang2024stableidentity, ). Due to the instability of GAN training, the generated embeddings have an unrestricted value range and fail to naturally insert into diffusion models under the setting of “w/o AdaIN” (decent results), as shown in Table 6. Since each dimension of word embedding has a different distribution, the AdaIN is introduced to contribute to the distribution alignment. Specifically, this operation helps the generated pseudo identity embeddings land more naturally into the celeb embedding space, so that the generator can deceive the discriminator more easily. Therefore, the introduced AdaIN will further unleash the prompt-aligned identity-consistent generation ability of IDE-GAN.
@@ Appendix C Limitations
Although the proposed CharacterFactory enables end-to-end identity-consistent character generation, it is not free of limitations. First, GAN-based methods can generate high-quality images, but these images may look unnatural with artifacts (karras2019style, ; fu2022stylegan, ). Similarly, our IDE-GAN may produce suboptimal embeddings, lacking consistency. Second, since the UNet of Stable Diffusion (SD) is fixed, we inherit not only the powerful generation ability, but also its drawbacks, such as hand anomalies (zhang2023detecting, ).

### v2

@@ 4.3. Comparison with Consistent-T2I Methods
In addition, we compare the most related methods The Chosen One (Avrahami et al., 2023) and ConsiStory (Tewel et al., 2024) with the content provided in their papers. These two methods are also designed for consistent character generation, but have not released the codes yet.
@@ 4.5. Interpolation Property of IDE-GAN
The interpolation property of GANs is that interpolations between different randomly sampled latent codes in latent space can produce semantically smooth variations in image space (Sainburg et al., 2018). To evaluate whether our IDE-GAN carries this property, we randomly sample $z_{1}$ and $z_{2}$, and perform linear interpolation as shown in Figure 6. IDE-GAN uses the interpolated latent codes to generate corresponding pseudo identity embeddings, respectively. Since the output space of IDE-GAN is embeddings instead of images, it cannot directly visualize the variations like traditional GANs (Sainburg et al., 2018; Karras et al., 2019) in image space. So we insert these pseudo identity embeddings into Stable Diffusion to generate the corresponding images via the pipeline in Figure 2(b). As shown in Figure 6, CharacterFactory can produce continuous identity variations with the interpolations between different latent codes. And the interpolated latent codes (e.g., $0.5z_{1}+0.5z_{2}$) can be chosen for further identity-consistent generation. It demonstrates that our IDE-GAN has satisfactory interpolation property and can be seamlessly integrated with Stable Diffusion.
@@ 4.6. Applications
Virtual Humans in Image/Video/3D Generation. Currently, virtual human generation mainly includes 2D/3D facial reconstruction, talking-head generation and body/human movements (Zhen et al., 2023), which typically rely on pre-existing images and lack scenario diversity and editability. And CharacterFactory can create new characters end-to-end and conduct identity-consistent virtual human image generation. In addition, since the pretrained Stable Diffusion 2.1 is fixed and the generated pseudo identity embeddings can be inserted into CLIP text transformer naturally, our method can collaborate with the SD-based plug-and-play modules. As shown in Figure 1, we integrate CharacterFactory with ControlNet-OpenPose (Zhang et al., 2023b; Cao et al., 2017), ModelScopeT2V (Wang et al., 2023) and LucidDreamer (Liang et al., 2023) to implement identity-consistent virtual human image/video/3D generation.
Identity-Consistent Dateset Construction. Some human-centric subject-driven generation methods (Li et al., 2023; Chen et al., 2023a) construct large-scale celeb datasets for training. PhotoMaker (Li et al., 2023) crawls celeb photos from the Internet and DreamIdentity (Chen et al., 2023a) uses text prompts containing celeb names to drive Stable Diffusion to generate celeb images. Their constructed data includes only celebs, leading to a limited number of identities. Notably, the proposed CharacterFactory can use diverse text prompts to generate identity-consistent images infinitely for dataset construction. Furthermore, collaboration with the mentioned SD-based plug-and-play modules can construct identity-consistent video/3D datasets.
@@ Appendix A More Results
In addition, we provide identity-consistent Image/Video/3D generation results as shown in Figure 9. Seamless integration with ControlNet (Zhang et al., 2023b), ModelScopeT2V (Wang et al., 2023) and LucidDreamer (Liang et al., 2023) empowers identity-consistent virtual human image/video/3D generation.
@@ Appendix B More Implementation Details
The Dimension of Latent Code $z$. Similar to traditional GAN models (Karras et al., 2019; Fu et al., 2022), the input of our Identity-Embedding GAN (IDE-GAN) is randomly sampled from a latent space, i.e., $z\in\mathcal{N}(0,\mathbf{I})$. Differently, our objective is to learn a mapping from the latent space to the celeb embedding space, i.e., $G:z\rightarrow v$. The CLIP word embedding’s dimension of Stable Diffusion 2.1 is 1024, so the dimension of IDE-GAN’s output $[v_{1}^{*},v_{2}^{*}]$ is $[2,1024]$. Based on this, we conduct experiments for the dimension of $z$, as shown in Table 4. Overall, IDE-GAN can generate consistent pseudo identity conforming to the given text prompts with different dimensions of $z$. However, we empirically find that a larger dimension of $z$ introduces more trainable parameters, leading to harder convergence and degraded identity consistency. A smaller $z$ represents narrower input space, resulting in poor identity diversity (this metric is designed in Section 4.4 for the diversity of generated identities). Therefore, we choose the setting of “Dim(z)=64” to trade off identity consistency and identity diversity, which performs well across the board.
@@
More Details of IDE-GAN. IDE-GAN is made up of MLPs, where the generator is $64-2048-2048$ MLP architecture and the discriminator is $2048-512-256-1$ MLP architecture. In addition to the ablation experiments in Section 4.4 of the main paper, we further explore the effect of AdaIN (Wang et al., 2024b). Due to the instability of GAN training, the generated embeddings have an unrestricted value range and fail to naturally insert into diffusion models under the setting of “w/o AdaIN” (decent results), as shown in Table 5. Since each dimension of word embedding has a different distribution, the AdaIN is introduced to contribute to the distribution alignment. Specifically, this operation helps the generated pseudo identity embeddings land more naturally into the celeb embedding space, so that the generator can deceive the discriminator more easily. Therefore, the introduced AdaIN will further unleash the prompt-aligned identity-consistent generation ability of IDE-GAN.
@@ Appendix C Limitations
Although the proposed CharacterFactory enables end-to-end identity-consistent character generation, it is not free of limitations. First, GAN-based methods can generate high-quality images, but these images may look unnatural with artifacts (Karras et al., 2019; Fu et al., 2022). Similarly, our IDE-GAN may produce suboptimal embeddings, lacking consistency. Second, since the UNet of Stable Diffusion (SD) is fixed, we inherit not only the powerful generation ability, but also its drawbacks, such as hand anomalies (Zhang et al., 2023a).